OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (0, 3)?

OpenStudy (callisto):

"Which of the following equations" means there are choices, right?

OpenStudy (anonymous):

ya

OpenStudy (callisto):

For a parabola having vertex at (h,k), it should be like this: y = a(x-h)^2+k If not, then expand it/make the equation into this form.

OpenStudy (anonymous):

idk tbh

OpenStudy (anonymous):

y=\[\left( x -3 \right)^{2}\]

OpenStudy (anonymous):

@goformit100 will u help me

OpenStudy (goformit100):

No, sorry Sir

OpenStudy (anonymous):

im a girl and okay

OpenStudy (goformit100):

Sorry madam... i am just 19 years old

OpenStudy (anonymous):

o cool message me

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

"For a parabola having vertex at (h,k), it should be like this: y = a(x-h)^2+k" Can you identify what h and k are?

OpenStudy (anonymous):

I posted what I thought the answer is am I right

OpenStudy (callisto):

It is not correct.

OpenStudy (anonymous):

okay will u explain it

OpenStudy (anonymous):

y = x^2- 3 is this the answer @Callisto

OpenStudy (callisto):

Since we have (0,3) as the vertex. Compared with (h,k), we can identify that h=0, k=3 Put (h,k) into y=a(x-k)^2+k, What would you get?

OpenStudy (anonymous):

Y=x^2+3

OpenStudy (callisto):

That's it.